Solve system of equations by the addition method
5x=2+7y
3x+4y=34

Answer by ReadingBoosters(3246)

Starting with
5x=2+7y; rearrange as 5x-7y=2
3x+4y=34
Multiply the first equation by 3 and the second equation by -5 to solve for y.
3(5x-7y=2)
-5(3x+4y=34)
to get
15x - 21y = 6
-15x - 20y = -170
Add the equations
15x + -15x - 21y + - 20y = 6 + -170
Combine like terms
-41y = -164
Divide by -41
y = 4
...
Starting with
5x-7y=2
3x+4y=34
Multiply the first equation by 4 and the second equation by 7 to solve for x.
4(5x-7y=2)
7(3x+4y=34)
to get
20x - 28y = 8
21x + 28y = 238
Add the equations
20x + 21x - 28y + 28y = 8 + 238
41x = 246
x = 6
...
Check
5(6)=30; 2+7(4)=30
3(6)+4(4)=34
..........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves
www.MyHomeworkAnswers.com